# Equivariant $\mathcal{D}$-stability for Actions of Tensor Categories

### v1

@@ Abstract.
We introduce a notion of equivariant $\mathcal{D}$-stability for actions of unitary tensor categories on $*{}^{*}$-algebras. We show that, when $\mathcal{D}$ is strongly self-absorbing, equivariant $\mathcal{D}$-stability of an action is equivalent to a unital embedding of $\mathcal{D}$ into a certain subalgebra of Kirchberg’s central sequence algebra. We use this to show $\mathcal{Z}$-stability for a large class of AF-actions.
SE was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 427320536 – SFB 1442, as well as under Germany’s Excellence Strategy EXC 2044 390685587, Mathematics Münster: Dynamics–Geometry–Structure, and was also partially supported by ERC Advanced Grant 834267 - AMAREC; SGP was partially supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by project G085020N funded by the Research Foundation Flanders (FWO); CJ was partially supported by NSF Grants DMS-2100531 and DMS-2247202
@@
Tensorial absorption of certain $*{}^{*}$-algebras plays a crucial role in the classification of simple amenable $*{}^{*}$-algebras and the investigation of their structure. In the infinite setting, tensorial absorption of the Cuntz algebra $\mathcal{O}_{\infty}$, a property known as $\mathcal{O}_{\infty}$-stability, powers the Kirchberg–Phillips theorem ([24, 32, 26]). In the finite setting, tensorial absorption of the Jiang–Su algebra $\mathcal{Z}$ has proven to be of fundamental importance ([18, 38, 14, 7, 37, 3, 2]).
Abstracting the key properties of the UHF algebras $\mathbb{M}_{n^{\infty}}$, the Cuntz algebras $\mathcal{O}_{2}$ and $\mathcal{O}_{\infty}$, and the Jiang–Su algebra $\mathcal{Z}$, Toms and Winter introduced the notion of a strongly self-absorbing $*{}^{*}$-algebra $\mathcal{D}$ and developed a unified theory of $\mathcal{D}$-stable $*{}^{*}$-algebras including machinery for detecting $\mathcal{D}$-stability in examples ([39]). The notion of $\mathcal{D}$-stability has subsequently been extended to the setting of group actions on $*{}^{*}$-algebras and has been successfully used both for the classification of group actions on $*{}^{*}$-algebras ([16, 35, 11, 17]) and to identify $\mathcal{Z}$-stability of crossed product $*{}^{*}$-algebras ([42, 12, 29, 30]).
@@
Employing the intertwining techniques developed in [13], we shall prove that equivariant $\mathcal{D}$-stability of an action $\mathcal{C}\curvearrowright A$ on a $*{}^{*}$-algebra $A$ is equivalent to the existence of a unital embedding of $\mathcal{D}$ into an appropriate subalgebra of the Kirchberg central sequence algebra of $A$ (Theorem 4.4). This result generalises [39, Theorem 2.2] and its group equivariant counterpart [36, Corollary 3.8] under the assumption that the acting group is discrete. We apply this characterisation to large class of AF-actions of unitary tensor categories (in the sense of [5]) and provide an easily verifiable criterion for determining equivariant $\mathcal{D}$-stability (Theorem 5.2). In particular, we show that these AF-actions are always equivariantly $\mathcal{Z}$-stable (Corollary 5.3).
@@
Our solution to this problem is to directly introduce a $*{}^{*}$-subalgebra of the central sequence algebra of $A$ that plays the role of the algebra of central sequences fixed under the action of the tensor category. The appropriate $*{}^{*}$-subalgebra is obtained by considering bounded sequences in $A$ whose the left and right actions asymptotically agree for all the Hilbert $A$-bimodules associated to the action (see Proposition 3.2). Since the tensor unit of $\mathcal{C}$ is mapped to the trivial $A$-bimodule, all such sequences are necessarily central sequences.
@@ 2. Preliminaries
We assume the reader is familiar with the language of tensor categories and $*{}^{*}$-tensor categories (see for example [8] and [13, Section 1.3 and Section 2]). We also assume some familiarity with both the theory of Hilbert modules (see for example [27]) and with strongly self-absorbing $*{}^{*}$-algebra as introduced in [39].
Let $A$ be a $*{}^{*}$-algebra. An $A$-$A$-correspondence is a $*{}^{*}$-homomorphism $\alpha:A\rightarrow\mathcal{L}(X)$ into the adjointable operators of a right Hilbert $A$-module $X$. This endows $X$ with the structure of a Hilbert $A$-bimodule in the sense of [22].
For $x\in X$ and $a\in A$ we denote the left action of $a$ on $x$ by $a\rhd x$ and the right action of $a$ on $x$ by $x\lhd a$. We denote the internal tensor product of Hilbert bimodules by $\boxtimes$. We say $\alpha$ is non-degenerate if $X=\overline{A\rhd X}=\overline{X\lhd A}$.
Let $\mathrm{Corr}_{0}(A)$ be the $*{}^{*}$-tensor category of non-degenerate Hilbert $A$-$A$-correspondences under the flipped internal tensor product of the bimodules, i.e. the monoidal product of $A$-$A$-correspondences $\alpha:A\rightarrow\mathcal{L}(X)$ and $\beta:A\rightarrow\mathcal{L}(Y)$ is the $A$-$A$-correspondence given by the internal tensor product $Y\boxtimes X$ (see [13, Section 1.2]).
@@
An action of a $*{}^{*}$-tensor category $\mathcal{C}$ on $A$ is a $*{}^{*}$-tensor functor $(G,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$.
An important example of a $*{}^{*}$-tensor category is the category of finite dimensional $\Gamma$-graded Hilbert spaces, where $\Gamma$ is a discrete group. We denote this category by $\mathrm{Hilb}(\Gamma)$. We write $\mathbb{C}_{\gamma}$ for $\Gamma$ graded Hilbert space that is $\mathbb{C}$ in the $\gamma$-th grading and zero elsewhere.
@@
We now recall the theory of dual objects in $*{}^{*}$-tensor categories as introduced in [28]. A $*{}^{*}$-tensor category $\mathcal{C}$ is said to be rigid if every $X\in\mathcal{C}$ admits a dual object. That is, there exists $\overline{X}\in\mathcal{C}$ and maps $R:1_{\mathcal{C}}\rightarrow\overline{X}\otimes X$ and $\overline{R}:1_{\mathcal{C}}\rightarrow X\otimes\overline{X}$ such that the zig-zag relations hold. Surpressing the associator and unitor maps these relations are
@@
Any rigid $*{}^{*}$-tensor category that admits subobjects is semisimple and hence every object can be decomposed as a finite direct sum of simple objects. When $\mathcal{C}$ is semisimple, we denote by $\mathrm{Irr}(\mathcal{C})$ a choice of isomorphism classes for simple objects.
@@
A unitary tensor category $\mathcal{C}$ is a rigid, semisimple, $*{}^{*}$-tensor category with simple unit object, i.e. $\mathrm{End}(1_{\mathcal{C}})\cong\mathbb{C}$. Moreover, $\mathcal{C}$ is called a unitary fusion category if it has finitely many isomorphism classes of simple objects.
In [22] Kajiwara, Pinzari and Watatani characterise which objects in $\mathrm{Corr}_{0}(A)$ have duals. It is shown in [22, Theorem 4.13] that if $E$ is a non-degenerate Hilbert $A$-bimodule admitting a dual object, then $E$ admits a compatible left $A$-valued inner product making $E$ a bi-Hilbertian $*{}^{*}$-bimodule in the sense of [22, Definition 2.3]. In this case, its dual is isomorphic to the contragredient bimodule $\overline{E}$.
@@
Notions of equivariant morphisms between two actions $(G,J)$ and $(H,I)$ of a $*{}^{*}$-tensor category $\mathcal{C}$ on a $*{}^{*}$-algebras $A$ and $B$ have appeared in the literature under (slightly) different names (see [5, 1, 13] for example). We shall use a definition of cocycle morphisms given by [5, Lemma 3.8] in the unital setting and by [13, Lemma 3.11] in general.
@@
A cocycle morphism $(\varphi,\{h^{X}\}_{X\in\mathcal{C}}):(G,J)\rightarrow(H,I)$ consists of a $*{}^{*}$-homomorphism $\varphi:A\rightarrow B$ and linear maps $h^{X}:G(X)\rightarrow H(X)$ for each $X\in\mathcal{C}$ such that
@@
Let $A$ be separable and $\mathcal{C}$ be semisimple with countably many isomorphism classes of simple objects. Cocycle morphisms $(\varphi,\{h^{X}\}_{X\in\mathcal{C}})$ and $(\psi,\{l^{X}\}_{X\in\mathcal{C}})$ from $(G,J)$ to $(H,I)$ are said to be approximately unitarily equivalent, if there exists a sequence of unitaries in the multiplier algebra $(u_{n})_{n\in\mathbb{N}}\in M(B)$ such that
@@ 3. A subalgebra of Kirchberg’s central sequences
Let $A$ be a $*{}^{*}$-algebra. We set $A_{\infty}=\ell_{\infty}(A)/c_{0}(A)$. Given a Hilbert $A$-$A$ bimodule $E$, we endow $E_{\infty}=\ell_{\infty}(E)/c_{0}(E)$ with the natural Hilbert $A_{\infty}$-bimodule structure and we identify $E$ with the image of constant sequences in $E_{\infty}$.
For any $*{}^{*}$-subalgebra $B\subset A_{\infty}$ the relative commutant and annihilator are given by
@@
Note that if $E$ is the trivial $A$-$A$ bimodule this agrees with usual definition of $A_{\infty}\cap A^{\prime}$. In general, $A_{\infty}\cap E^{\prime}$ is a norm closed subalgebra of $A_{\infty}$. If we also consider the dual bimodule $\overline{E}$, we get a $*{}^{*}$-subalgebra.
@@
Let $A$ be a $*{}^{*}$-algebra and $E$ be a dualisable Hilbert $A$-bimodule. Then the norm closed algebra $(A_{\infty}\cap E^{\prime})\cap(A_{\infty}\cap\overline{E}^{\prime})$ is a $*{}^{*}$-algebra.
@@
Let $B=(A_{\infty}\cap E^{\prime})\cap(A_{\infty}\cap\overline{E}^{\prime})$, as the intersection of two closed subalgebras of $A_{\infty}$, it follows that $B$ is also a closed subalgebra of $A_{\infty}$. It remains to show that $B$ is a $*{}^{*}$-subalgebra of $A_{\infty}$. Let $a\in B$ and $\eta\in E$ then
@@
is a $*{}^{*}$-algebra. Moreover, if $A$ is $\sigma$-unital $F(A)^{G}$ is unital.
@@
As $\mathcal{C}$ is rigid, any $X\in\mathcal{C}$ has a dual object $(\overline{X},R,\overline{R})$. So each correspondence $G(X)$ also has a dual object in $\mathrm{Corr}_{0}(A)$ given by $(G(\overline{X}),J_{\overline{X},X}^{*}G(R),J_{X,\overline{X}}^{*}G(\overline{R}))$. Therefore, $G(\overline{X})\cong\overline{G(X)}$ and it follows from Lemma 3.1 that $\bigcap_{X\in\mathcal{C}}A_{\infty}\cap G(X)^{\prime}$ is a $*{}^{*}$-subalgebra of $A_{\infty}\cap A^{\prime}$.
@@
If $A$ is unital the annihilator $A_{\infty}\cap A^{\perp}$ is trivial and we simply denote the $*{}^{*}$-algebra of Proposition 3.2 by $(A_{\infty}\cap A^{\prime})^{G}$.
@@
Let $\Gamma$ be a discrete group and $\mathcal{C}=\mathrm{Hilb}(\Gamma)$ be acting by automorphisms on a $*{}^{*}$-algebra $A$. That is, the $\mathrm{Hilb}(\Gamma)$ action is given by a cocycle action $(\alpha,u)$ and sends $\mathbb{C}_{\gamma}\mapsto{}_{\alpha_{\gamma}}A$ for $\gamma\in\Gamma$. Then $a\in\bigcap_{\gamma\in\Gamma}A_{\infty}\cap({}_{\alpha_{\gamma}}A)^{\prime}$ if and only if $(\alpha_{\gamma}(a)-a)b=0$ for all $b\in A$. Therefore, $F(A)^{(\alpha,u)}$ coincides with the fixed point algebra $F(A)^{\alpha}$ of those central sequences fixed by the automorphisms $\alpha_{\gamma}$ for all $\gamma\in\Gamma$ modulo $A_{\infty}\cap A^{\perp}$.

### v2

@@ Abstract.
We introduce a notion of equivariant $\mathcal{D}$-stability for actions of unitary tensor categories on C-algebras. We show that, when $\mathcal{D}$ is strongly self-absorbing, equivariant $\mathcal{D}$-stability of an action is equivalent to a unital embedding of $\mathcal{D}$ into a certain subalgebra of Kirchberg’s central sequence algebra. We use this to show $\mathcal{Z}$-stability for a large class of AF-actions.
SE was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 427320536 – SFB 1442, as well as under Germany’s Excellence Strategy EXC 2044 390685587, Mathematics Münster: Dynamics–Geometry–Structure, and was also partially supported by ERC Advanced Grant 834267 - AMAREC; SGP was partially supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by projects G085020N and 1249225N funded by the Research Foundation Flanders (FWO); CJ was partially supported by NSF Grants DMS-2100531 and DMS-2247202
@@
Tensorial absorption of certain C-algebras plays a crucial role in the classification of simple amenable C-algebras and the investigation of their structure. In the infinite setting, tensorial absorption of the Cuntz algebra $\mathcal{O}_{\infty}$, a property known as $\mathcal{O}_{\infty}$-stability, powers the Kirchberg–Phillips theorem ([24, 32, 26]). In the finite setting, tensorial absorption of the Jiang–Su algebra $\mathcal{Z}$ has proven to be of fundamental importance ([18, 38, 14, 7, 37, 3, 2]).
Abstracting the key properties of the UHF algebras $\mathbb{M}_{n^{\infty}}$, the Cuntz algebras $\mathcal{O}_{2}$ and $\mathcal{O}_{\infty}$, and the Jiang–Su algebra $\mathcal{Z}$, Toms and Winter introduced the notion of a strongly self-absorbing C-algebra $\mathcal{D}$ and developed a unified theory of $\mathcal{D}$-stable C-algebras including machinery for detecting $\mathcal{D}$-stability in examples ([39]). The notion of $\mathcal{D}$-stability has subsequently been extended to the setting of group actions on C-algebras and has been successfully used both for the classification of group actions on C-algebras ([16, 35, 11, 17]) and to identify $\mathcal{Z}$-stability of crossed product C-algebras ([42, 12, 29, 30]).
@@
Employing the intertwining techniques developed in [13], we shall prove that equivariant $\mathcal{D}$-stability of an action $\mathcal{C}\curvearrowright A$ on a C-algebra $A$ is equivalent to the existence of a unital embedding of $\mathcal{D}$ into an appropriate subalgebra of the Kirchberg central sequence algebra of $A$ (Theorem 4.4). This result generalises [39, Theorem 2.2] and its group equivariant counterpart [36, Corollary 3.8] under the assumption that the acting group is discrete. We apply this characterisation to large class of AF-actions of unitary tensor categories (in the sense of [5]) and provide an easily verifiable criterion for determining equivariant $\mathcal{D}$-stability (Theorem 5.2). In particular, we show that these AF-actions are always equivariantly $\mathcal{Z}$-stable (Corollary 5.3).
@@
Our solution to this problem is to directly introduce a C-subalgebra of the central sequence algebra of $A$ that plays the role of the algebra of central sequences fixed under the action of the tensor category. The appropriate C-subalgebra is obtained by considering bounded sequences in $A$ whose the left and right actions asymptotically agree for all the Hilbert $A$-bimodules associated to the action (see Proposition 3.2). Since the tensor unit of $\mathcal{C}$ is mapped to the trivial $A$-bimodule, all such sequences are necessarily central sequences.
@@ 2. Preliminaries
We assume the reader is familiar with the language of tensor categories and C-tensor categories (see for example [8] and [13, Section 1.3 and Section 2]). We also assume some familiarity with both the theory of Hilbert modules (see for example [27]) and with strongly self-absorbing C-algebra as introduced in [39].
Let $A$ be a C-algebra. An $A$-$A$-correspondence is a -homomorphism $\alpha:A\rightarrow\mathcal{L}(X)$ into the adjointable operators of a right Hilbert $A$-module $X$. This endows $X$ with the structure of a Hilbert $A$-bimodule in the sense of [22].
For $x\in X$ and $a\in A$ we denote the left action of $a$ on $x$ by $a\rhd x$ and the right action of $a$ on $x$ by $x\lhd a$. We denote the internal tensor product of Hilbert bimodules by $\boxtimes$. We say $\alpha$ is non-degenerate if $X=\overline{A\rhd X}=\overline{X\lhd A}$, which for $A$ unital occurs precisely when the left and right actions are unital.
Let $\mathrm{Corr}_{0}(A)$ be the C-tensor category of non-degenerate Hilbert $A$-$A$-correspondences under the flipped internal tensor product of the bimodules, i.e. the monoidal product of $A$-$A$-correspondences $\alpha:A\rightarrow\mathcal{L}(X)$ and $\beta:A\rightarrow\mathcal{L}(Y)$ is the $A$-$A$-correspondence given by the internal tensor product $Y\boxtimes X$ (see [13, Section 1.2]).
@@
An action of a C-tensor category $\mathcal{C}$ on $A$ is a C-tensor functor $(G,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$.
An important example of a C-tensor category is the category of finite dimensional $\Gamma$-graded Hilbert spaces, where $\Gamma$ is a discrete group. We denote this category by $\mathrm{Hilb}(\Gamma)$. We write $\mathbb{C}_{\gamma}$ for $\Gamma$ graded Hilbert space that is $\mathbb{C}$ in the $\gamma$-th grading and zero elsewhere.
@@
We now recall the theory of dual objects in C-tensor categories as introduced in [28]. A C-tensor category $\mathcal{C}$ is said to be rigid if every $X\in\mathcal{C}$ admits a dual object. That is, there exists $\overline{X}\in\mathcal{C}$ and maps $R:1_{\mathcal{C}}\rightarrow\overline{X}\otimes X$ and $\overline{R}:1_{\mathcal{C}}\rightarrow X\otimes\overline{X}$ such that the zig-zag relations hold. Surpressing the associator and unitor maps these relations are
@@
Any rigid C-tensor category that admits subobjects is semisimple and hence every object can be decomposed as a finite direct sum of simple objects. When $\mathcal{C}$ is semisimple, we denote by $\mathrm{Irr}(\mathcal{C})$ a choice of isomorphism classes for simple objects.
@@
A unitary tensor category $\mathcal{C}$ is a rigid, semisimple, C-tensor category with simple unit object, i.e. $\mathrm{End}(1_{\mathcal{C}})\cong\mathbb{C}$. Moreover, $\mathcal{C}$ is called a unitary fusion category if it has finitely many isomorphism classes of simple objects.
In [22] Kajiwara, Pinzari and Watatani characterise which objects in $\mathrm{Corr}_{0}(A)$ have duals. It is shown in [22, Theorem 4.13] that if $E$ is a non-degenerate Hilbert $A$-bimodule admitting a dual object, then $E$ admits a compatible left $A$-valued inner product making $E$ a bi-Hilbertian C-bimodule in the sense of [22, Definition 2.3]. In this case, its dual is isomorphic to the contragredient bimodule $\overline{E}$.
@@
Notions of equivariant morphisms between two actions $(G,J)$ and $(H,I)$ of a C-tensor category $\mathcal{C}$ on a C-algebras $A$ and $B$ have appeared in the literature under (slightly) different names (see [5, 1, 13] for example). We shall use a definition of cocycle morphisms given by [5, Lemma 3.8] in the unital setting and by [13, Lemma 3.11] in general.
@@
A cocycle morphism $(\varphi,\{h^{X}\}_{X\in\mathcal{C}}):(G,J)\rightarrow(H,I)$ consists of a -homomorphism $\varphi:A\rightarrow B$ and linear maps $h^{X}:G(X)\rightarrow H(X)$ for each $X\in\mathcal{C}$ such that
@@
Let $A$ be a separable C-algebra and $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Cocycle morphisms $(\varphi,\{h^{X}\}_{X\in\mathcal{C}})$ and $(\psi,\{l^{X}\}_{X\in\mathcal{C}})$ from $(G,J)$ to $(H,I)$ are said to be approximately unitarily equivalent, if there exists a sequence of unitaries in the multiplier algebra $(u_{n})_{n\in\mathbb{N}}\in M(B)$ such that
@@ 3. A subalgebra of Kirchberg’s central sequences
Let $A$ be a C-algebra. We set $A_{\infty}=\ell_{\infty}(A)/c_{0}(A)$. Given a Hilbert $A$-$A$ bimodule $E$, we endow $E_{\infty}=\ell_{\infty}(E)/c_{0}(E)$ with the natural Hilbert $A_{\infty}$-bimodule structure and we identify $E$ with the image of constant sequences in $E_{\infty}$.
For any C-subalgebra $B\subset A_{\infty}$ the relative commutant and annihilator are given by
@@
Note that if $E$ is the trivial $A$-$A$ bimodule this agrees with usual definition of $A_{\infty}\cap A^{\prime}$. In general, $A_{\infty}\cap E^{\prime}$ is a norm closed subalgebra of $A_{\infty}$. If we also consider the dual bimodule $\overline{E}$, we get a C-subalgebra.
@@
Let $A$ be a C-algebra and $E$ be a dualisable Hilbert $A$-bimodule. Then the norm closed algebra $(A_{\infty}\cap E^{\prime})\cap(A_{\infty}\cap\overline{E}^{\prime})$ is a C-algebra.
@@
Let $B=(A_{\infty}\cap E^{\prime})\cap(A_{\infty}\cap\overline{E}^{\prime})$, as the intersection of two closed subalgebras of $A_{\infty}$, it follows that $B$ is also a closed subalgebra of $A_{\infty}$. It remains to show that $B$ is a -subalgebra of $A_{\infty}$. Let $a\in B$ and $\eta\in E$ then
@@
is a C-algebra. Moreover, if $A$ is $\sigma$-unital $F(A)^{G}$ is unital.
@@
As $\mathcal{C}$ is rigid, any $X\in\mathcal{C}$ has a dual object $(\overline{X},R,\overline{R})$. So each correspondence $G(X)$ also has a dual object in $\mathrm{Corr}_{0}(A)$ given by $(G(\overline{X}),J_{\overline{X},X}^{*}G(R),J_{X,\overline{X}}^{*}G(\overline{R}))$. Therefore, $G(\overline{X})\cong\overline{G(X)}$ and it follows from Lemma 3.1 that $\bigcap_{X\in\mathcal{C}}A_{\infty}\cap G(X)^{\prime}$ is a C-subalgebra of $A_{\infty}\cap A^{\prime}$.
@@
If $A$ is unital the annihilator $A_{\infty}\cap A^{\perp}$ is trivial and we simply denote the C-algebra of Proposition 3.2 by $(A_{\infty}\cap A^{\prime})^{G}$.
@@
Let $\Gamma$ be a discrete group and $\mathcal{C}=\mathrm{Hilb}(\Gamma)$ be acting by automorphisms on a C-algebra $A$. That is, the $\mathrm{Hilb}(\Gamma)$ action is given by a cocycle action $(\alpha,u)$ and sends $\mathbb{C}_{\gamma}\mapsto{}_{\alpha_{\gamma}}A$ for $\gamma\in\Gamma$. Then $a\in\bigcap_{\gamma\in\Gamma}A_{\infty}\cap({}_{\alpha_{\gamma}}A)^{\prime}$ if and only if $(\alpha_{\gamma}(a)-a)b=0$ for all $b\in A$. Therefore, $F(A)^{(\alpha,u)}$ coincides with the fixed point algebra $F(A)^{\alpha}$ of those central sequences fixed by the automorphisms $\alpha_{\gamma}$ for all $\gamma\in\Gamma$ modulo $A_{\infty}\cap A^{\perp}$.